OpenStudy (anonymous):

Will Medal and Fan What is the solution to the system of equations? 5x-3y=-9 2x-5y=4

OpenStudy (anonymous):

@One098 @Opcode @wio @TheAsker2002

OpenStudy (anonymous):

@xo.minnie.xox

OpenStudy (anonymous):

Can you isolate a variable in the first equation?

OpenStudy (anonymous):

What I mean is, for: \[ 5x-3y=9 \]Can you solve for \(x\) or \(y\)? Your first step is to solve for one of them.

OpenStudy (anonymous):

5x-3y=9 x=3 and y=2

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Well, that would be one solution, but we don't want to find a single solution for it, because it might not be the solution for the other equation.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

Suppose we want to isolate \(x\). First we would add \(3y\) to both sides: \[5x = 9+3y\]Then, we would divide by \(5\): \[ x= \frac{9}{5}+\frac {3y}{5}\]

OpenStudy (anonymous):

We can substitute this for the \(x\) in the other equations:\[ 2x-5y=4 \]Will become: \[ 2\left(\frac 95+\frac {3y}5\right)-5y=4 \]Now we can solve for \(y\).

OpenStudy (anonymous):

\[ \frac{18}5+\frac{6y}5-5y = 4 \]We should multiply everything by \(5\) to get rid of fractions: \[ 18+6y-25y=20 \]

OpenStudy (anonymous):

18+6y-25y=20 so every number in this equation multiplied by 5

OpenStudy (anonymous):

Hold on one second...

OpenStudy (anonymous):

I made a mistake, I was using\[ 5x-3y=\color{red}9 \]But I should be using \[ 5x-3y=\color{red}{-9} \]So, fixing that mistake results in: \[ -18+6y-25y=4 \]

OpenStudy (anonymous):

Now, we can combine the like terms for \(y\) to get: \[ -18-19y = 20 \]

OpenStudy (anonymous):

ohhh ok is that our last step to the equation?

OpenStudy (anonymous):

Then we add \(18\): \[ -19y = 38 \]And this gives us \(y=-2\). Does that make sense?

OpenStudy (anonymous):

Now, to solve \(x\), we substitute our solution for \(y\).\[ 5x-3(-2)=-9 \]

OpenStudy (anonymous):

Can you solve for \(x\)?

OpenStudy (anonymous):

yes thank you for not giving me the answer and explaning :)

OpenStudy (anonymous):

@wio how does that go to my answer choices A. (0, 4) B. (–4, 0) C. (0, –4) D. (4, 0)

OpenStudy (anonymous):

Sorry, but those answer choices can't be for the same question.

OpenStudy (anonymous):

Use the graph to state the solution for the system. x – y = 4 (line a) x + y = 4 (line b) A. (0, 4) B. (–4, 0) C. (0, –4) D. (4, 0)

OpenStudy (anonymous):

The answer is \((x,y) = (-3,-2)\).

OpenStudy (anonymous):

Well, this is a separate question, but the point where they intersect is the solution.

OpenStudy (anonymous):

And they appear to intersect at \((4,0)\).

OpenStudy (anonymous):

Thank you for ur time! : )

OpenStudy (anonymous):

im gonna open another question can you help me with that one to @wio

OpenStudy (anonymous):

Sorry, but I have to go soon.